# MK-SGN: A Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation for Skeleton-based Action Recognition

### v4

@@ 3 Proposed Method
This section begins with the comprehensive architectural design of the MK-SGN, followed by an introduction to its Spiking Coding mechanism, the Spiking Multimodal Fusion Module, SGN Block and the Knowledge Distillation Method from GCN to SGN are then described in detail.
@@ 3.1 Overall Architecture
Our network architecture, illustrated in Figure 2, comprises a 10-layer multimodal teacher GCN network and a 6-layer SNN student network. The teacher GCN network processes the four multimodal skeleton datasets separately. The processed features from each modality are then passed through the Global Average Pooling (GAP) and Fully Connected (FC) layers to generate predictions. These predictions are finally fused to produce the soft labels.
The student SGN network processes each modality independently through the SSC module, transforming skeleton data into spike-form representations. Spike-form features are then fused via the SMF module, utilizing mutual information to optimize cross-modal integration. Each student network layer includes an SA-SGC module and a STC module, which are designed to capture and refine temporal and spatial features efficiently and effectively. Our approach implements knowledge distillation through inter-layer features and soft label distillation. The inter-layer features from the GCN teacher network are converted into spike-form representations via the FTM, ensuring alignment with the spiking domain and facilitating the training of the SNN student network. Soft labels produced by the teacher network provide supplementary supervisory signals, optimizing the learning process. The subsequent sections provide a detailed description of the design and functionality of the modules.
@@ 3.2 Spiking Coding Mechanism and Spiking Multimodal Fusion
This chapter discusses how to convert continuous real skeleton data into spiking-form representation. Then, we describe how to effectively fuse the multimodal spike-form features derived from multimodal skeleton data.
@@
(b) The detail structure of the SMIC Module
@@ 3.2.1 Skeleton Spiking Coding Module
For a given 3D skeleton sequence data $X_{i}\in\mathbb{R}^{C\times T\times V}$, the SSC Module first transforms it into an image-like representation. Then it expands it into a form suitable for spiking temporal representation illustrated in Figure 2(a). The process is described as follows:
$P_{0}=\operatorname{Expand}(X_{i})\in\mathbb{R}^{S\times C\times V\times T},$ (3)
$P_{i}=\operatorname{SN}(\operatorname{BN}(\operatorname{Conv2d}(P_{0})))\in\mathbb{R}^{S\times N\times V\times T}.$ (4)
In the initial step, the skeleton sequence data is transformed into an image-like representation to facilitate the structured feature extraction of spatial and temporal characteristics in subsequent stages. Subsequently, the $\operatorname{Expand}(\cdot)$ operation adds a spike time step dimension to the data, reorganizing the temporal structure to align with the requirements of spiking representation. The resulting expanded representation $P_{0}$ incorporates the additional spike time step dimension $S$ while preserving the original temporal dimension $T$, the graph structure $V$, and the channel size $C$.
[…]
@@ 3.2.2 Spike-based Multimodal Fusion Module
Multimodal skeleton data consists of four components: bones ($X_{\mathrm{B}}$), joints ($X_{\mathrm{J}}$), bone motions ($X_{\mathrm{BM}}$), and joint motions ($X_{\mathrm{JM}}$), each represented with dimensions $(C,T,V)$.
The SMF module is a mutual information-based fusion framework formulated specifically for the sparse spike-form features derived from multimodal data. In the SMF, the multimodal skeleton data are first encoded independently through the SSC module to obtain their respective spike-form features: $P_{\mathrm{B}},P_{\mathrm{J}},P_{\mathrm{BM}},P_{\mathrm{JM}}\in\mathbb{R}^{S\times D\times V\times T}$.
The spike-form features are then pairwise selected and concatenated along the feature dimension $D$ to construct a joint distribution. To construct a corresponding marginal distribution, one of the two spike-form features in each pair is randomly shuffled along the spike time step dimension $S$, and the same concatenation is performed. We propose the Spike-form Mutual Information Calculation (SMIC) network, formally constituted by an LSTM network, SN, FC, and GAP layers, to realize the spiking-driven mutual information prediction. The framework computes the mutual information between four sparse spike-form features and performs weighted fusion. By leveraging the computational efficiency of sparse spike-form features and maximizing the lower bound of mutual information, the SMIC network effectively captures shared information across modalities, facilitating both robust and energy-efficient multimodal feature fusion within SNN framework, as illustrated in Figure 3(b).
@@
Both $P_{\mathrm{Joint}}$ and $P_{\text{marginal}}$ are then processed through an LSTM network equipped with SN layers. For the joint distribution, the output of the LSTM is passed through a FC layer followed by a GAP operation to compute:
@@
Using the predictions $\boldsymbol{t}$ and $\boldsymbol{et}$, the lower bound of mutual information is estimated as:
$mi_{-}lb=\bar{\boldsymbol{t}}-\log\overline{\boldsymbol{et}}.$ (9)
@@
$\boldsymbol{\overline{MI}}=\frac{\sum_{i=1}^{4}\boldsymbol{MI}_{i,:}}{\sum_{i=1}^{4}\sum_{j=1}^{4}\boldsymbol{MI}_{i,j}},$ (11)
$\boldsymbol{W}=\frac{\boldsymbol{\overline{MI}}-\min(\boldsymbol{\overline{MI}})}{\max(\boldsymbol{\overline{MI}})-\min(\boldsymbol{\overline{MI}})}\in\mathbb{R}^{4}.$ (12)
The fused spike-form features are then obtained by weighting the original features of all modalities using $\boldsymbol{W}$:
$P_{\text{fuse}}=\boldsymbol{W}*\left[P_{\mathrm{B}},P_{\mathrm{J}},P_{\mathrm{BM}},P_{\mathrm{JM}}\right]^{\boldsymbol{T}}\in\mathbb{R}^{S\times D\times V\times T}.$ (13)
Finally, the fused feature $P_{\text{fuse}}$ is fed into SGN Block for subsequent processing. By leveraging spiking-driven mutual information, the SMF module ensures the effective integration of sparse multimodal features, enhancing the robustness and accuracy of the overall action recognition framework.
@@ 3.3 SGN Block
Revisiting vanilla skeleton-based action recognition based on GCN, we construct our SGN using features converted into spike-form and adjacency matrices. The process can be summarized as follows:
$S=\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}},\quad$ (14)
where $S$ is the normalized adjacency matrix, $\widetilde{A}=A+I$ denotes the sum of the adjacency matrix $A$ and the identity matrix $I$ (representing self-connections). $\widetilde{D}$ is the diagonal degree matrix of $\widetilde{A}$. The graph convolutional output is then calculated using the following formula:
$H=S^{K}P\in\mathbb{R}^{S\times N\times T\times V},$ (15)
where, $H$ represents the output of the early spatial spiking graph convolution, $K$ is the number of graph convolutional layers, and $P$ is the spike-form feature input. The first layer of SGC can be expressed as:
$H^{(0)}=\mathrm{SN}(P)+\mathrm{SN}(HW_{sgc}^{(0)}),\quad$ (16)
where $W_{sgc}^{(0)}\in\mathbb{R}^{N^{(0)}\times N^{(1)}}$ is the learnable weight matrix of the first layer. Each subsequent layer has its learnable weights, denoted as $W_{sgc}^{(l)}\in\mathbb{R}^{N^{(l)}\times N^{(l+1)}}$.
@@
To further enhance the model’s representation capabilities, we incorporate Self Spiking Attention (SSA) following the principles of existing spiking attention methods. These methods leverage a spiking-based $QKV$ structure with scaling factors $s$ to control large matrix multiplication values. The computation of SSA can be described as follows:
$Q^{(0)}=\mathrm{SN}_{Q}\left(\mathrm{BN}\left(H_{0}W^{(0)}_{Q}\right)\right),\quad$ (17)
$K^{(0)}=\mathrm{SN}_{K}\left(\mathrm{BN}\left(H_{0}W^{(0)}_{K}\right)\right),\quad$ (18)
$V^{(0)}=\mathrm{SN}_{V}\left(\mathrm{BN}\left(H_{0}W^{(0)}_{V}\right)\right),\quad$ (19)
$H^{(0)}_{SA}=H_{0}+\operatorname{SN}\left((Q^{(0)}{K^{(0)}}^{T})V^{(0)}\cdot s\right),\quad$ (20)
@@
where, $Q^{(0)}$, $K^{(0)}$, and $V^{(0)}$ are computed through spiking neurons ($\mathrm{SN}$) after Batch Normalization ($\mathrm{BN}$) and linear transformations using learnable weights $W^{(0)}_{Q},W^{(0)}_{K},$ and $W^{(0)}_{V}$. The attention mechanism integrates the weighted interactions, while the scaling factor $s$ prevents large values in the spiking attention computation. $H^{(0)}_{SA}$ combines the residual connection $H_{0}$ with the attention-enhanced feature, defined as SA-SGC, illustrated in Figure 2(b), yielding a robust representation for spiking graph processing. After that, the first layer of STC can be written as:
$T^{(0)}=\operatorname{SN}\left(\mathrm{BN}\left(\operatorname{Conv2d}\left(H^{(0)}_{SA}\right)\right)\right)+\operatorname{SN}\left(H^{(0)}_{SA}\right).$ (21)
@@
For subsequent layers, the spiking self attention-enhanced feature $H^{(l)}_{SA}$ is updated and combined with temporal features $T^{l-1}$, expressed as,illustrated in Figure 2(c):
$H^{(l)}_{SA}=\operatorname{SN}\left(T^{(l-1)}\right)+\operatorname{SN}\left(H^{(l-1)}_{SA}W_{sgc}^{(l-1)}\right),$ (22)
$T^{(l)}=\operatorname{SN}\left(\mathrm{BN}^{(l)}\left(\operatorname{Conv2d}^{(l)}\left(H^{(l)}_{SA}\right)\right)\right)+\operatorname{SN}\left(H^{(l)}_{SA}\right).$ (23)
@@
Finally, the processed features are sent through GAP and a FC layer to output the prediction, as follows:
@@
where $y\in\mathbb{R}^{U}$, and $U$ is the number of action classes. Based on this, a standard cross-entropy loss is adopted for model optimization, formulated as:
$L_{\mathrm{task}}=-\frac{1}{B}\sum_{b=1}^{B}\sum_{u=1}^{U}\mathrm{Y}_{bu}\log\mathrm{y}_{bu}+\left(1-\mathrm{Y}_{bu}\right)\log\left(1-\mathrm{y}_{bu}\right),$ (25)
where, $\mathrm{Y}$ is the ground-truth label, $\mathrm{y}$ is the predicted output, and $B$ denotes the number of training samples.
@@ 3.4 Knowledge Distillation of GCN-to-SGN
The proposed approach fully exploits the rich knowledge embedded in pre-trained multimodal GCNs, transferring superior representational features to the SGN. By aligning the intermediate features and outputs of the SGN with those of the GCN, shown in Figure 4(a), the method significantly enhances SGN accuracy. The enhancement is achieved without introducing additional computational complexity or energy consumption, thereby preserving the inherent efficiency advantages of the SGN architecture. The distillation framework bridges the performance gap between GCNs and SGNs while maintaining the suitability of SGNs for energy-efficient applications.
@@ 3.4.1 Soft Label Distillation
In the proposed distillation framework, the CTR-GCN with a 4-stream architecture is utilized as the teacher network. Multimodal data inputs, including bone, joint, bone-motion, and joint-motion data, are processed independently by single-stream CTR-GCN. Single-modal soft labels, represented as $y_{b}$, $y_{j}$, $y_{bm}$, and $y_{jm}$, are generated corresponding to the respective modalities.
The outputs of the four modalities are then aggregated to form the multimodal soft labels $\hat{y}$, which serve as the knowledge to be distilled into the SGN. This aggregation is performed using weighted summation, as described below:
$\hat{y}=\alpha_{b}y_{b}+\alpha_{j}y_{j}+\alpha_{bm}y_{bm}+\alpha_{jm}y_{jm},$ (26)
where $\alpha_{b}$, $\alpha_{j}$, $\alpha_{bm}$, and $\alpha_{jm}$ are hyperparameters that control the relative contribution of each modality. These parameters are set based on the importance or reliability of the modalities during training.
The SGN then learns to approximate the multimodal teacher soft labels $\hat{y}$ through a regression-based loss function. The final learning objective for soft label distillation is expressed as:
$L_{sdk}=\|y-\hat{y}\|,$ (27)

### v5

@@ 3 Proposed Method
This section provides an overview of MK-SGN’s comprehensive architecture, followed by a detailed explanation of its Skeleton Spiking Coding module and Spiking Multimodal Fusion Module. Additionally, the SGN Block and the Knowledge Distillation Method from GCN to SGN are thoroughly discussed.
Figure 2: The overview of the proposed MK-SGN architecture. The teacher network comprises a 10-layer GCN that processes four multimodal skeleton data separately, followed by Global Average Pooling (GAP) and Fully Connected (FC) layers to generate soft labels. The student network is a 6-layer SGN that incorporates the following key modules: (a) The Skeleton Spiking Coding (SSC) module, responsible for transforming raw skeleton data into spike-form feature to ensure compatibility with spiking neural computations. (b) The Spiking Self-Attention Spiking Graph Convolution (SA-SGC) module is formulated to model spatial dependencies and enhance feature extraction by integrating Spiking Graph Convolution and Spiking Self-Attention (SSA). (c) The Spiking Temporal Convolution (STC) module is proposed to capture temporal patterns and refine temporal dynamics within spike-form features. Additionally, the Spike-based Multimodal Fusion Module (SMF) optimizes cross-modal integration based on mutual information, and the Feature Transformation Module (FTM) maps inter-layer GCN features into spike-form features for effective alignment with the SGN. Knowledge distillation is achieved through the FTM and supervision from the soft labels produced by the teacher network.
@@ 3.1 Overall Architecture
Our network architecture, illustrated in Figure 2, comprises a 10-layer multimodal teacher GCN modal and a 6-layer student SGN modal. The teacher GCN processes the four multimodal skeleton data separately. The processed features from each modality are then passed through the Global Average Pooling (GAP) and Fully Connected (FC) layers to generate predictions. These predictions are finally fused to produce the soft labels. The student SGN network processes each modality through the SSC module, transforming skeleton data into spike-form representations. Spike-form features are then fused via the SMF module, which utilizes mutual information to optimize cross-modal integration. Each student network layer includes the SA-SGC and STC modules, which are formulated to capture and refine spatial and temporal features efficiently and effectively. Our approach implements knowledge distillation through inter-layer features and soft label distillation. The inter-layer features from the teacher GCN modal are converted into spike-form representations via the FTM, which ensures compatibility with the spiking neural domain and facilitates the training of the student SGN modal. Soft labels produced by the teacher modal provide supplementary supervisory signals, optimizing the learning process. The subsequent sections provide a detailed account of the formulation of the modules.
@@ 3.2 Spiking Coding Mechanism and Spiking Multimodal Fusion
This chapter explores the conversion from continuous real-valued skeleton data into spiking-form representations. It then details the methods for fusing multimodal spike-form feature derived from multimodal skeleton data.
@@
(b) The detail structure of the Spike-form Mutual Information Calculation (SMIC) Module
@@ 3.2.1 Skeleton Spiking Coding Module
For a given 3D skeleton sequence data $X_{\mathrm{i}}\in\mathbb{R}^{C\times T\times V}$, the SSC Module first transforms it into an image-like representation. Then, it expands it into a form suitable for spiking temporal representation illustrated in Figure 2(a). The process is described as follows:
$P_{0}=\operatorname{Expand}(X_{\text{i}})\in\mathbb{R}^{S\times C\times V\times T},$ (3)
$P_{\text{i}}=\operatorname{SN}(\operatorname{BN}(\operatorname{Conv2d}(P_{0})))\in\mathbb{R}^{S\times D\times V\times T}.$ (4)
In the initial step, the skeleton sequence data is transformed into an image-like representation to facilitate the structured feature extraction of spatial and temporal characteristics in subsequent stages. Subsequently, the $\operatorname{Expand}(\cdot)$ operation adds a spike time step dimension $S$ to the data, reorganizing the temporal structure to align with the requirements of spiking representation. The expanded representation $P_{0}$ incorporates the additional spike time step dimension $S$ while preserving the original temporal dimension $T$, the graph structure $V$, and the channel size $C$.
Following expansion, the representation $P_{0}$ is processed through a 2D convolutional layer ($\operatorname{Conv2d}$) with a kernel size of 3, followed by Batch Normalization ($\operatorname{BN}$) and a Spiking Neuron ($\operatorname{SN}$) layer. The pipeline refines the temporal-spatial features within the spiking domain, yielding a spiking feature representation $P$ with dimensions $S\times D\times V\times T$. The spike time step dimension $S$ reflects the temporal structure of spiking activity, while the dimension $T$ retains the temporal characteristics of the original sequence, ensuring a comprehensive temporal-spatial representation. $D$ represents the hidden channel size.
@@ 3.2.2 Spike-based Multimodal Fusion Module
Multimodal skeleton data consists of four components: bones ($X_{\mathrm{b}}$), joints ($X_{\mathrm{j}}$), bone motions ($X_{\mathrm{bm}}$), and joint motions ($X_{\mathrm{jm}}$), each represented with dimensions $(C,T,V)$. The multimodal skeleton data are first represented independently through the SSC module to obtain their respective spike-form features: $P_{\mathrm{b}},P_{\mathrm{j}},P_{\mathrm{bm}},P_{\mathrm{jm}}\in\mathbb{R}^{S\times D\times V\times T}$ and then fed into the SMF module. The SMF module is based on mutual information for the fusion of sparse spike-form features derived from multimodal data.
The spike-form features are then pairwise selected and concatenated along the feature dimension $D$ to construct a joint distribution. To construct a corresponding marginal distribution, one of the two spike-form features in each pair is randomly shuffled along the spike time step dimension $S$, and the same concatenation is performed. We propose the Spike-form Mutual Information Calculation (SMIC) module, formally composed of LSTM, SN, FC, and GAP layers, to realize spiking-driven mutual information prediction. The framework computes the mutual information between four sparse spike-form features and performs weighted fusion. By leveraging the computational efficiency of sparse spike-form features and maximizing the lower bound of mutual information, the SMIC network effectively captures shared information across modalities, facilitating robust and energy-efficient multimodal feature fusion within SNN framework, as illustrated in Figure 3(b).
@@
Both $P_{\mathrm{Joint}}$ and $P_{\text{marginal}}$ are then processed through an LSTM network equipped with SN layers. For the joint distribution, the output of the LSTM is passed through a FC layer followed by a GAP operation as:
@@
Based on the predictions $\boldsymbol{t}$ and $\boldsymbol{et}$, the lower bound of mutual information is estimated as:
$mi\_lb_{\mathrm{jb}}=\bar{\boldsymbol{t}}-\log\overline{\boldsymbol{et}}.$ (9)
@@
$\boldsymbol{\overline{MI}}=\frac{\sum_{r=1}^{4}\boldsymbol{MI}_{r,:}}{\sum_{r=1}^{4}\sum_{c=1}^{4}\boldsymbol{MI}_{r,c}},$ (11)
$\boldsymbol{W}_{\text{MI}}=\frac{\boldsymbol{\overline{MI}}-\min(\boldsymbol{\overline{MI}})}{\max(\boldsymbol{\overline{MI}})-\min(\boldsymbol{\overline{MI}})}\in\mathbb{R}^{4}.$ (12)
The fused spike-form features are then obtained by weighting the original features of all modalities using $\boldsymbol{W}_{\text{MI}}$:
$P_{\text{fuse}}=\boldsymbol{W}_{\text{MI}}*\left[P_{\mathrm{b}},P_{\mathrm{j}},P_{\mathrm{bm}},P_{\mathrm{jm}}\right]^{\boldsymbol{T}}\in\mathbb{R}^{S\times D\times V\times T}.$ (13)
Finally, the fused feature $P_{\text{fuse}}$ is fed into the SGN Block for subsequent processing. By leveraging spiking-driven mutual information, the SMF module ensures the effective integration of sparse multimodal features, enhancing the robustness and accuracy of the overall framework.
@@ 3.3 SGN Block
Revisiting vanilla skeleton-based action recognition based on GCN, we construct our SGN with spike-form features and adjacency matrices. The process can be summarized as follows:
$\mathcal{A}=\Delta^{-\frac{1}{2}}\widetilde{A}\Delta^{-\frac{1}{2}},\quad$ (14)
where $\mathcal{A}$ is the normalized adjacency matrix of dimension $\times V\times V. KK$ represents the number of branches, each capturing distinct relationships between nodes, and $V$ denotes the number of nodes in the graph. $\widetilde{A}=A+I$ denotes the sum of the adjacency matrix $A$ and the identity matrix $I$ (representing self-connections), while $\Delta$ is the diagonal degree matrix of $\widetilde{A}$.
To calculate the multi-branch graph convolution, the branch-specific adjacency matrix $\mathcal{A}_{k}\in\mathbb{R}^{V\times V}$ is applied to the input spiking-form feature $P\in\mathbb{R}^{S\times D\times T\times V}$. The first layer of the SGC begins by aggregating outputs from all $K$ branches and is defined as:
$H^{(0)}=\mathrm{SN}\left(\mathrm{BN}\left(PW_{\text{r}}^{(0)}\right)\right)+\mathrm{SN}\left(\mathrm{BN}\left(\sum_{k=1}^{K}\mathcal{A}_{k}PW_{k}^{(0)}\right)\right),$ (15)
@@
where $\mathcal{A}_{k}$ represents the adjacency matrix for branch $k$, $W_{k}^{(0)}\text{ and }W_{\text{r}}\in\mathbb{R}^{D\times D^{\prime}}$ are the learnable weight matrix specific to the $k$-th branch and a residual connection, and $\mathrm{SN}$ applied to the aggregated features. To further enhance the model’s representation capabilities, we incorporate SSA following the principles of existing spiking attention methods[36]. These methods leverage a spiking-based $QKV$ structure with scaling factors $s$ to control large matrix multiplication values. The computation of SSA can be described as follows:
$Q^{(0)}=\mathrm{SN}\left(\mathrm{BN}\left(H^{(0)}W^{(0)}_{Q}\right)\right)$$\quad$ (16)
$K^{(0)}=\mathrm{SN}\left(\mathrm{BN}\left(H^{(0)}W^{(0)}_{K}\right)\right)$$\quad$ (17)
$V^{(0)}=\mathrm{SN}\left(\mathrm{BN}\left(H^{(0)}W^{(0)}_{V}\right)\right)$$\quad$ (18)
$H^{(0)}_{\text{SA}}=H^{(0)}+\operatorname{SN}\left((Q^{(0)}{K^{(0)}}^{T})V^{(0)}\cdot s\right),\quad$ (19)
@@
where, $Q^{(0)}$, $K^{(0)}$, and $V^{(0)}$ are computed through $\mathrm{SN}$ after $\mathrm{BN}$ and linear transformations using learnable weights $W^{(0)}_{Q},W^{(0)}_{K}$, and $W^{(0)}_{V}$. The attention mechanism integrates the weighted interactions, while the scaling factor $s$ prevents large values in the spiking attention computation. $H^{(0)}_{\text{SA}}$ combines the residual connection $H^{(0)}$ with the attention-enhanced feature, defined as SA-SGC, illustrated in Figure 2(b). SA-SGC yields a robust representation for spiking graph processing. After that, the first layer of STC, illustrated in Figure 2(c), can be written as:
$T^{(0)}=\operatorname{SN}\left(\mathrm{BN}\left(W_{\text{t}}^{(0)}\left(H^{(0)}_{\text{SA}}\right)\right)\right)+\operatorname{SN}\left(H^{(0)}_{\text{SA}}\right)$, (20)
@@
where $W_{\text{t}}$ is the learnable weight matrix. For subsequent layers, the spiking self attention-enhanced feature $H^{(l)}_{\text{SA}}$ and temporal features $T^{(l-1)}$, expressed as:
$H^{(l)}=\mathrm{SN}\left(\mathrm{BN}\left(T^{(l-1)}W_{\text{r}}^{(l)}\right)\right)+\mathrm{SN}\left(\mathrm{BN}\left(\sum_{k=1}^{K}\mathcal{A}_{k}T^{(l-1)}W_{k}^{(l)}\right)\right),$ (21)
$H^{(l)}_{\text{SA}}=H^{(l)}+\operatorname{SN}\left((Q^{(l)}{K^{(l)}}^{T})V^{(l)}\cdot s\right),\quad$ (22)
$T^{(l)}=\operatorname{SN}\left(\mathrm{BN}^{(l)}\left(W_{\text{t}}^{(l)}\left(H^{(l)}_{\text{SA}}\right)\right)\right)+\operatorname{SN}\left(H^{(l)}_{\text{SA}}\right)$. (23)
@@
Finally, the output features of the ast layer $L$ are processed through GAP and a FC layer to generate the prediction, as follows:
@@
where $y\in\mathbb{R}^{U}$, and $U$ is the number of action classes. We further calculate the standard cross-entropy loss $\mathcal{L}_{\mathrm{task}}$ based on the predicted probabilities and one-hot encoded ground-truth labels.
@@ 3.4 Knowledge Distillation of GCN-to-SGN
The proposed approach fully exploits the rich knowledge embedded in pre-trained multimodal GCNs, transferring superior representational features to the SGN. By aligning the intermediate features and outputs of the SGN with those of the GCN, shown in Figure 4(a), the method significantly enhances SGN accuracy. The enhancement is achieved without introducing additional computational complexity or energy consumption during the inference phase, thereby preserving the inherent efficiency advantages of the SGN architecture. The distillation framework bridges the performance gap between GCNs and SGNs while maintaining the suitability of SGNs for energy-efficient applications.
@@ 3.4.1 Soft Label Distillation
In the proposed distillation framework, the CTR-GCN with a 4-stream architecture is utilized as the teacher network [68]. Multimodal data inputs, including bone, joint, bone-motion, and joint-motion data, are processed independently by single-stream CTR-GCN. Single-modal soft labels, represented as $y_{\text{b}}$, $y_{\text{j}}$, $y_{\text{bm}}$, and $y_{\text{jm}}$, are generated corresponding to the respective modalities.
The outputs of the four modalities are then aggregated to form the multimodal soft labels $y_{\text{mm}}$, which serve as the knowledge to be distilled into the SGN. This aggregation is performed using weighted summation, as described below:
$y_{\text{mm}}=\alpha_{\text{b}}y_{\text{b}}+\alpha_{\text{j}}y_{\text{j}}+\alpha_{\text{bm}}y_{\text{bm}}+\alpha_{\text{jm}}y_{\text{jm}},$ (25)
where $\alpha_{\text{b}}$, $\alpha_{\text{j}}$, $\alpha_{\text{bm}}$, and $\alpha_{\text{jm}}$ are hyperparameters that control the relative contribution of each modality. These parameters are set based on the importance or reliability of the modalities during training.
The SGN then learns to approximate the multimodal teacher soft labels $y_{\text{mm}}$ through a regression-based loss function. The final learning objective for soft label distillation is expressed as:
$\mathcal{L}_{\text{sdk}}=\|y-y_{\text{mm}}\|,$ (26)